Thursday, January 21, 2010

Journal Page: January 21

It’s finally clear, and not very cold – about -3C. Been cloudy for weeks, and suddenly a good clear stretch. Quite warm – around freezing – the last few days so humidity will be high and probably poor transparency.

Tonight’s plan is to gather more data for the B33 region that I started a couple of weeks ago. I still need all three RGB colours, and would also like some straight luminance to serve as a base for the Ha capture I already have. I will also do a quick RGB image of a G2V star, if I can find a convenient one quickly, to calibrate the colour mix.

Last time I wasted evening by taking all calibration frames first and then having to stop. Tonight I’ll gather data and take calibration at the end. Or not at all – might be able to reuse from 2 weeks ago since I haven’ moved the camera.

Plan for tonight, to be modified no doubt:

  • Take 30 minutes each of R, G, and B on B33 area, binned 2×2, using 6×5-minute subs
    • Do the colours separately, focusing with each filter
      • Red
      • Green
      • Blue
  • Take 90 minutes of luminance, unbinned, using 10-minute subs
    • Luminance
  • Take darks
  • Take flats and bias
  • Take RGB of a G2V star in the general area of Orion.
    • HIP 24336, HIP 29248, HIP 23831 are near Orion
    • HIP 29248 looks best – nicely separated. RA 6-10-39 Dec +17-55-56 (now)
      • It’s GSC 1318:1060, which Maxim can find.

Set up and framed. Weird way-off in goto calibration but got it. framed b33.
autoguider wouldn’t calibrate – cable problem. get spare, and longer, cables – I think stress from stretched cables pulled connector out of connection.

Red filtered 6×5-min. Not much tme available – already gettingn near next-door house roof. might get 3 colours.

PEC on but getting wide oscillation in Y. Turned Y aggressiveness up to 7, seems to have controlled it a bit. Then up to 7.5. Figure it’s right when it’s oscillating across zero on each sample (like x is doing)

Accumulated 25 minutes of Red successfully. During 6th exposure, lost guide star behind next door roof. End of session. -9 degrees by the time I finished.

Tried switching to M1 but scope pointing was completely screwed – I think it hit the slew limits while I was inside. Will need to re-do sky model again because I had to manually return it to CWD position. Need to start earlier to get G, B, and L data.

Dropped laptop out of case onto deck while putting away – broke corner plastic. Seems OK.

Processing: Can’t use the calibration frames from 2 weeks ago with Maxim and these images – says “no applicable frames”. Don’t know why. Next time, get new calibration frames and try cycling colours so I get at least some. Also do Meridian flip so I don’t hit RA stop during exposure.

Leave a Reply

Your email address will not be published. Required fields are marked *

This site uses Akismet to reduce spam. Learn how your comment data is processed.